the digits 1, 2, 3, 4, 5, and 7 are randomly arranged to form a six-digit number. Find the probability that the first and last digits are both even.
Answer by sudhanshu_kmr(1152)


no. of ways to arrange these 6 digits = 6!

no. of ways to arrange 6 digits where first and last digit are even
= no. of ways to arrange first and last digit * no. of ways to arrange 4 inner digits
= 2! * 4!
= 2 * 24
= 48

required probability = 48/6!
= 48/720
= 1/15